OpenStudy (anonymous):

Solve the system of linear equations. 6x + 3y = 33 4x + y = 15

OpenStudy (anonymous):

6x + 3y = 33 4x + y = 15<---changes to y=-4x+15 Substitute. 6x + 3(-4x + 15) = 33 6x -12x +45 = 33 Combine like terms -6x + 45 = 33 Subtract 45 on both sides -6x = -12 Divide by -6 on both sides x = 2 Plug x back into one of the equations. 6(2) + 3y = 33 12 + 3y = 33 Subtract 12 on both sides 3y = 21 Divide by 3 on both sides y = 7 (2,7) is your final answer.